OpenStudy (anonymous):

ASAP!!!!! I need how the problem is set up, the work solving the problem, along with the solution of them: 1. A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?

OpenStudy (anonymous):

Let the speed of the airplane in still air be x. On the 5 hr trip from orlando to denver, the speed of the plane is x+ 40. Note that tailwind means wind blowing in the direction of where you are travelling. Distance travelled = (x+40) * 5 On the 6 hr trip from denver to orlando, the speed of the plane x-40. Distance travelled = (x-40) * 6 Since both distance travelled is equal, (x+40) *5 = (x-40) * 6 5x+ 200 = 6x - 240 x = 440. Speed of the airplane 440 miles per hour